### v1

# Diffuson-Dominated Thermal Transport Crossover from Ordered to Liquid-like Cu3BiS3:The Negligible Role of Ion Hopping

Abstract
Fundamentally understanding lattice dynamics and thermal transport behavior in liquid-like, partially occupied compounds remains a long-standing challenge in condensed matter physics. Here, we investigate the microscopic mechanisms underlying the ultralow thermal conductivity in ordered/liquid-like Cu3BiS3 by combining experimental methods with first-principles calculations. We first experimentally synthesize and characterize the ordered structure and liquid-like, partially Cu-atom occupied Cu3BiS3 structure with increasing temperature. We then combine self-consistent phonon calculations, including bubble-diagram corrections, with the Wigner transport equation, considering both phonon propagation and diffuson contributions, to evaluate the anharmonic lattice dynamics and thermal conductivity in phase-change Cu3BiS3. Our theoretical model predicts an ultralow thermal conductivity of 0.34 W·m-1·K-1 at 400 K, dominated by diffuson contributions, which accurately reproduces and explains the experimental data. Importantly, the machine-learning-based molecular dynamics (MD) simulations not only reproduced the partially Cu-atom occupied Cu3BiS3 structure with the space group but also successfully replicated the thermal conductivity obtained from experiments and Wigner transport calculations. This observation highlights the negligible impact of ionic mobility arising from partially occupied Cu sites on the thermal conductivity in diffuson-dominated thermal transport compounds. Our work not only sheds light on the minimal impact of ionic mobility on ultralow thermal conductivity in phase-change materials but also demonstrates that the Wigner transport equation accurately describes thermal transport behavior in partially occupied phases with diffuson-dominant thermal transport.
keywords:
American Chemical Society, LaTeXThese authors contributed equally
[…]
Thayer School of Engineering, Dartmouth College, Hanover, New Hampshire, 03755, USA
[…]
Linköping University, Department of Physics, Chemistry, and Biology (IFM), Linköping, SE-58183, Sweden
[…]
State Key Laboratory of Superhard Materials, College of Physics, Jilin University, Changchun, 130012, China
Keywords: thermoelectric material, Cu-based chalcogenide, lattice anharmonicity, wigner transport equation, self-consistent phonon theory
[…]
Some journals require a graphical entry for the Table of Contents. This should be laid out “print ready” so that the sizing of the text is correct.
Inside the tocentry environment, the font used is Helvetica 8 pt, as required by Journal of the American Chemical Society.
The surrounding frame is 9 cm by 3.5 cm, which is the maximum permitted for Journal of the American Chemical Society graphical table of content entries. The box will not resize if the content is too big: instead it will overflow the edge of the box.
This box and the associated title will always be printed on a separate page at the end of the document.
1 Introduction
Ultra-low lattice thermal conductivity () in solid-state materials is crucial for advancing energy conversion technologies 1, 2, improving fuel cell performance 3, and optimizing various thermoelectric applications in the fields of thermal barrier coatings 4, waste heat recovery 5, 6. However, understanding and measuring lattice thermal conductivity is particularly challenging for materials undergoing phase transitions 7. These materials are widely available and have been extensively used in various fields, including thermoelectrics 8, solid-state memory 9, and switches 10. Therefore, a fundamental understanding of the origins of ultra-low thermal conductivity through a comprehensive investigation of the relationships between transport phenomena, lattice dynamics, and structural configurations is of great significance. This understanding unveils critical insights and establishes foundational principles for the innovation and discovery of advanced material systems 11.
Within the family of phase transition materials, Cu-based chalcogenides stand out due to their complex phase transition behaviors and outstanding thermoelectric properties 12, 13. Cu-based chalcogenides often undergo complex lattice distortions and significant changes in thermal transport properties during phase transitions, making them a key platform for exploring new thermal conduction mechanisms. For instance, Liu et al. 14 reported a significant reduction in the thermal conductivity of Cu2Se, attributed to its second-order phase transition, which led to a notable enhancement of the thermoelectric figure of merit () from 0.6 to 2.3. The recent synthesis of copper-rich wittichenite Cu3BiS3 15, 16, which exhibits intrinsically low lattice thermal conductivity, is particularly noteworthy. This material provides a compelling platform to investigate structural features—such as atomic coordination and cationic networks—that influence thermal transport. Meanwhile, the present investigation on Cu3BiS3 clearly demonstrates that the 3-fold coordination of copper, combined with its strong anisotropic vibrations, serves as the driving force behind its ultra-low thermal conductivity 17. Despite extensive research on Cu3BiS3, its phase transition behavior remains unclear, and the mechanisms underlying its ultra-low thermal conductivity across different phases have yet to be fully elucidated.
Historically, thermal transport in crystalline materials has been well described by the phonon-gas model, which treats phonons as quasiparticles propagating over long distances within the framework of the conventional phonon Boltzmann transport equation (BTE) 18, 19. Within the conventional BTE framework, point defects 20, anharmonicity 21, 22, and bond strength variations 23, 24 act as phonon scattering sources, reducing mean free paths and thereby lowering lattice thermal conductivity. However, certain complex crystalline materials exhibit ultra-low thermal conductivities that cannot be fully explained by the phonon-gas model 25. In such cases, phonon mean free paths are extremely short—comparable to or even smaller than interatomic distances—resulting in the loss of their quasi-particle-like nature 26. To address the limitations of the conventional BTE, Simoncelli et al. 27 developed a unified theory of thermal transport by reformulating the Boltzmann Transport Equation within the Wigner formalism. Specifically, the Wigner transport equation 28 describes thermal transport via the phonon velocity operator, wherein the diagonal and off-diagonal elements capture both the particle-like propagation of phonons and the wave-like tunneling between eigenstates. The wave-like tunneling arises from the overlap of phonon modes with similar frequencies—substantially broadened by anharmonicity—thereby facilitating thermal energy transfer between modes. The unified theory of thermal transport has been successfully applied to address the underestimation of thermal conductivity within the BTE framework for strongly anharmonic systems, such as Cu12SbS29, La2Zr2O727, Cs2AgBiBr6 30.
Another challenge in evaluating the thermal conductivity of phase-change materials is the potential occurrence of a partial-crystalline partial-liquid state (PCPLS), driven by the liquid-like migration of Cu ions, as observed in superionic systems such as Ag9GaSe631, AgCrSe232, and Li2S 33. It is also reported that the most representative -Cu2S, with space group and Pearson symbol cF208, exhibits a remarkably high degree of disorder, with 8 Cu ions distributed over 204 possible sites within the unit cell 34, 35. The strikingly structural disorder suggests an exceptionally low activation barrier for Cu-ion migration, giving rise to liquid-like cationic dynamics. This remarkable ion mobility facilitates strong phonon scattering mechanisms, thereby reducing the specific heat via the suppression of transverse phonon modes 36. In contrast, Bernges et al. 37 conducted the first direct comparison of thermal and ionic transport in nine Ag+ argyrodite compositions. Their results revealed that, despite the ionic conductivity varying by several orders of magnitude, the thermal conductivity remains largely constant, indicating that high ionic mobility is not essential for diffusion-mediated thermal transport. Similarly, Ghata et al. 13 observed that with increasing temperature, the ionic conductivity of Cu7PSe6 rises significantly, while the thermal conductivity remains nearly constant, suggesting a negligible influence of ionic transport on thermal conductivity. Therefore, a fundamental understanding of the thermal transport mechanism in phase-transition and partially occupied materials remains an open question and an urgent need.
In this work, we systematically investigate the anharmonic lattice dynamics and thermal transport behavior in ordered and elemental-partially occupied, liquid-like Cu3BiS3 by combining experimental and theoretical approaches. Experimentally, we synthesize and characterize the ordered and liquid-like phases of Cu3BiS3, which adopt the space groups and , respectively. We then combine first-principles calculations with machine-learning-based molecular dynamics simulations to accurately reproduce the partially Cu-atom occupied phase of Cu3BiS3 () at high temperatures. Subsequently, we applied the state-of-the-art Wigner transport equation, accounting for both propagative and diffuson contributions, to calculate the thermal conductivity of the ordered and phases of Cu3BiS3, achieving excellent agreement with experimental results. To investigate the impact of ion hopping on thermal transport at high temperatures in Cu3BiS3, we calculate the lattice thermal conductivity () using the machine-learning-based Green–Kubo equilibrium molecular dynamics (GK-EMD) method 38, which explicitly accounts for all contributions to the heat-flux operator 39, 40. We find that the predicted thermal conductivity accurately reproduces both the experimental data and the results obtained from the Wigner transport equation, indicating a negligible effect of ion mobility on thermal transport. Finally, we explore the microscopic origins of the ultralow thermal conductivity observed in both phases of Cu3BiS3. Our work not only highlights the negligible role of ionic mobility in thermal transport but also provides deeper insights into the microscopic mechanisms underlying the ultralow thermal conductivity observed in ordered and liquid-like Cu3BiS3 and related compounds.
2 Results and discussion
2.1 Experimental crystal structures
Rietveld refinement of the powder X-ray diffraction (PXRD) data, carried out using the GSAS-II software package 41, 42, shows excellent agreement with the calculated patterns (R = 7.86%, R = 10.22%, and = 1.12), as illustrated in Figure 1(a). As evidenced by the well-defined PXRD patterns at 300 K, the compound is accurately indexed to the non-centrosymmetric space group , consistent with previously reported results 17, 43. Specifically, the structure comprises three Cu atoms, three S atoms, and one Bi atom, all positioned at Wyckoff 4a sites with a site multiplicity of 4. Each Bi atom is bonded to three S atoms, forming a [BiS3]3- trigonal pyramid that contains a lone pair of electrons. In contrast, each Cu atom is coordinated by three S atoms, forming a distorted trigonal planar arrangement. The unit cell parameters determined from Rietveld refinement of the phase-pure sample, namely, a = 7.6750(2) Å, b = 10.4083(3) Å, and c = 6.7060(2) Å, are in good agreement with density functional theory (DFT) calculations performed using the PBEsol functional 44 (a = 7.4618(8) Å, b = 10.4705(9) Å, c = 6.3209(5) Å) as well as with previously reported results 16, 17.
Upon further increasing the temperature, we observe a notable change in the XRD pattern (see Figure S1 in the Supporting Information). Based on the structural refinments of the high-T PXRD pattern, we identify an experimental phase transition from the orthorhombic space group to the orthorhombic space group is observed, as illustrated in Figure 1(b). This finding diverges from the previously proposed space group for CuBiS2 reported by Wu et al.in Ref.16. Although our results differ from those reported by Wu et al., it is noteworthy that the isostructural space Cu3SbS3 has been experimentally observed to undergo a phase transition from the (-) phase to the (-) phase 45, 46. This observation aligns with our findings, supporting the validity of the observed phase transition in Cu3BiS3. Interestingly, the distribution of disordered Cu+ ions within the three-dimensional framework of [BiS3]3- units has been elucidated in the phase, matching that observed in -Cu3SbS3 47. Specifically, single-crystal data reveal that Cu+ ions exhibit disorder, being partially distributed over five distinct sites, resulting in either trigonal-planar or tetrahedral coordination. To investigate this phenomenon, we conducted 20 ps ab initio molecular dynamics (AIMD) simulations at 500 K (see Figure 1(c) and Figure S2 in the Supporting Information). Notably, the simulations successfully captured the diverse sites occupied by Cu+ ions, confirming their mobility and aligning closely with our experimental observations.
To elucidate the chemical environment of Cu+ ions, we performed electronic structure calculations on the phase, examining both bonding and antibonding states to gain deeper insights into its bonding characteristics. In the phase, various Cu–S bonds exhibit distinct bonding characteristics (see Figure 1(d)); nevertheless, these bonds generally share common features. All Cu–Se bonds exhibit occupied antibonding states below the Fermi level, contributing approximately 45% to the overall antibonding character. Previous studies have attributed this behavior to the proximity of the Cu d-orbitals to the energy levels of the S p-orbitals, leading to strong p–d hybridization 48. Such robust hybridization pushes the occupied antibonding states closer to the Fermi level, weakening the chemical bonds and promoting metavalent-like behavior 24, 49. The combination of low integrated electron population and substantial antibonding contributions highlights the intrinsically weak bonding environment of Cu+ within the structure, thus facilitating rapid Cu+ ion transport 50. Furthermore, the significant presence of occupied antibonding states below the Fermi level notably enhances lattice anharmonicity, substantially influencing thermal transport, as discussed in subsequent sections. Also, such a weak bonding leads to larger mean square displacement (MSD) of Cu atoms compared to the other atoms (see Figure S3 in the Supporting Information). This is consistent with the experimental observation of the atomic displacement parameters of Cu3BiS3 atoms 17.
2.2 Diffuson-dominanted thermal transport
To accurately capture the finite-temperature lattice dynamics, the anharmonic phonon frequencies of crystalline Cu3BiS3 are calculated using self-consistent phonon theory with bubble correction (SCPB) 30, 51 (see Figures 2(a) and 2(b)). Interestingly, from Figures 2(a) and 2(b), we find that in both the and Pnma phases, low-frequency phonons exhibit softening with increasing temperature, in significant contrast to previously reported results for other crystalline compounds, such as BaZrO3 26, TlInTe2 52, and YbFe4Sb12 53. Note that in strongly anharmonic materials, such as CsPbBr351, KNbO3and NaNbO354, the negative energy shift from the bubble diagram 51 is significant and imposes a non-nigligible impact on both the theoretical phase transition temperature and the phonon linewidth. Specifically, considering only the energy shift from quartic anharmonicity leads to low-frequency phonon hardening in both the and Pnma phases of Cu3BiS3 (see Figure S4 in the Supporting Information). Therefore, incorporating the full explicit anharmonic vibrational effects, including both cubic and quartic anharmonicity, enables a more accurate description of phonon softening, which significantly influences the lattice thermal conductivity.
With the finite-temperature phonons obtained, we proceed to calculate the lattice thermal conductivity of Cu3BiS3 using the Wigner transport equation 27, 28, accounting for both phonon population and coherence contributions, as shown in Figure 2(c). Our results show that the thermal conductivity of the phase at 400 K reaches 0.36 Wm-1K-1, in good agreement with the experimental value of 0.34 Wm-1K-1 17, thereby validating our calculations. More specifically, at 400 K, the coherence contribution to the total thermal conductivity is 0.22 Wm-1K-1, accounting for 6̃1% of the total thermal conductivity, highlighting the dominant role of phonon wave-like tunneling and strong lattice anharmonicity 55. We next continue to calculate the thermal conductivity of the ordered Pnma phase of Cu3BiS3 above 400 K using the Wigner transport equation 27, 28(see in Figure 2(c)). Our theoretical model predicts an ultra-low total thermal conductivity of 0.36 Wm-1K-1, comprising 0.16 Wm-1K-1 from the population contribution and 0.20 Wm-1K-1 from the coherence contribution. Surprisingly, the predicted thermal conductivity of the ordered Pnma phase of Cu3BiS3 agrees well with the experimentally observed value of 0.34 Wm-1K-1 at 400 K for the partially occupied Pnma phase. Furthermore, as temperature increases, the theoretical predictions remain in good agreement with experimental measurements for partial-occupied Cu3BiS3 (Pnma space group), as shown in Figure 2(c). The cross-phase thermal transport in Cu3BiS3 remains unexplored 16, motivating us to compare the predicted thermal conductivities of both the and Pnma phases with experimental values, as shown in the inset of Figure 2(c). Our results show that the thermal conductivity of the phase increases sharply above 400 K, deviating from the experimental values. In contrast, the Pnma phase shows a continuous decrease in thermal conductivity above 400 K, showing a good agreement with experimental results at high temperatures. These results further confirm the temperature-induced phase transition from to Pnma and provide a clear explanation for the smooth change in thermal conductivity at the transition point.
Considering the liquid-like behavior of the Pnma phase at high temperatures, we calculate using the Green-Kubo equilibrium molecular dynamics (GK-EMD) method 38 (see Figure S5 in the Supporting Information), which accounts for all terms in the heat flux operator 39, 40. As shown in Figure 2(c), the total thermal conductivity predicted by the GK-EMD method at 400 K is in excellent agreement with that obtained from the Wigner transport equation 27, 28. Even at 500 K, the discrepancy between the two methods remains negligible, with a difference of only 0.01 Wm-1K-1. Based on the above observation, and considering that the thermal conductivity from the Wigner transport equation 27, 28 is calculated for the ordered Pnma phase—i.e., in the absence of ion hopping—we conclude that ion hopping has minimal impact on thermal transport in liquid-like Cu3BiS3. The negligible effect of ion hopping on thermal transport can be attributed to the fact that ion-hopping behavior resembles rattling-like motion observed in certain crystalline materials 13, 56, which induces strong anharmonicity. This, in turn, suppresses phonon transport while enhancing the coherence contribution. Thus, in diffuson-dominated thermal transport regimes, ion hopping contributes minimally to the total thermal conductivity.
To further validate our theoretically predicted lattice thermal conductivity of crystalline Cu3BiS3, we compared the calculated values with experimental measurements in the temperature range of 2 to 100 K (see Figure 2(d)). Clearly, when boundary scattering at low temperatures is not considered 18, the predicted thermal conductivity—accounting only for three-phonon, four-phonon, and phonon-isotope scattering—deviates significantly from the experimental values. To accurately reproduce the experimental thermal conductivity 57, 58, effective grain sizes of 50, 100, and 200 nm are introduced, which significantly suppress the thermal conductivity and bring the predictions closer to the experimental results. On top of the 200 nm boundary scattering model, the inclusion of additional point-defect scattering leads to improved agreement between theory and experiment, as depicted in Figure 2(d). In addition, we performed a comparative analysis between the experimentally measured isobaric heat capacity (Cp) of Cu3BiS3 and the corresponding theoretical predictions within the temperature range of 2–70 K (see Figure S6 in the Supporting Information). Notably, at low temperatures, the experimental Cp values exhibit excellent consistency with the calculated constant volume heat capacity (Cv), reflecting the reliability of the theoretical model in this regime. Meanwhile, our computational results closely reproduce the isobaric heat capacity data reported by Wu et al. prior to the observed phase transition 16.
2.3 Microscopic mechanisms of thermal transport
Next, to reveal the microscopic mechanisms of thermal transport in crystalline Cu3BiS3 with the and phases, we calculate the spectral and cumulative propagating thermal conductivity at 300 K and 400 K, respectively. As shown in Figure 3(a), phonons with frequencies below 1.5 THz dominate the particle-like phonon channel, which can be attributed to the relatively large group velocities of the acoustic and low-frequency optical modes (see Figure S7 in the Supporting Information). To elucidate the correlation between the ultra-low propagating thermal conductivity and the atomic structure of crystalline Cu3BiS3, we project the atomic contributions onto the phonon bands, as shown in Figure 3(b). Specifically, Cu atoms predominantly contribute to the low- and mid-frequency optical phonon modes in the range of 1.5–5.0 THz. In contrast, S atoms primarily participate in the high-frequency optical modes above 5.0 THz, while Bi atoms exhibit relatively low participation and are mainly involved in a limited number of low-frequency optical modes between 1.5 and 2.0 THz (see Figure S8 in the Supporting Information). The fully Cu- and Bi-occupied low-frequency optical modes lead to strong phonon scattering rates, thereby suppressing thermal transport in crystalline Cu3BiS3, a phenomenon similarly observed in CsCu4Se3 55. Importantly, Cu atoms give rise to low-energy, Einstein-like quasilocalized vibrational modes 56, particularly within the 2–4 THz frequency range. These Cu-dominated quasilocalized vibrational phonon modes act as strong phonon scatterers, significantly interacting with and disrupting the propagation of heat-carrying acoustic phonons. This phenomenon has also been reported in cubic SrTiO359, the double perovskite Cs2AgBiBr630, and copper-based chalcogenides such as o-CsCu5S324.
As illustrated in Figure 3(c), we further calculate the phonon lifetimes for the phase at 300 K and the phase at 400 K, respectively. We find that the majority of phonon modes exhibit lifetimes shorter than the Wigner limit 27, 28, considering both three-phonon (3ph) and four-phonon (4ph) scattering processes, indicating strong anharmonicity in the system. For the phase, we observe a dip in phonon lifetimes for phonons around 0.75 THz, which are dominated by Bi atoms, highlighting a strong scattering source that contributes to the suppression of thermal conductivity 26, 30. Similarly, the flattened, localized vibrational mode, primarily associated with Cu atoms near 2.2 THz, also exhibits an extremely short lifetime and pronounced anharmonic scattering (see Figure 3(c) and Figure S9 in the Supporting Information). Therefore, we further confirm that the ultra-low thermal conductivity is attributed to the vibrations of Cu and Bi atoms. In the phase, Cu atoms play a similar role, as the Cu-dominated mode near 1.8 THz exhibits a significant drop in phonon lifetime (see Figure 3(c) and Figure S9 in the Supporting Information), corresponding to the dip observed in the spectral thermal conductivity (see Figure 3(a)). Although strong anharmonicity and the low phonon lifetimes of quasilocalized modes suppress particle-like phonon propagation 25, they enhance the wave-like phonon diffuson contribution (see Figure 3(d)). Theoretically, the presence of dense, quasilocalized phonon modes with similar energies and strong anharmonicity can facilitate significant heat exchange between phonon modes in energy space 60, 55, thereby promoting diffusion-like transport 61.
To gain a better understanding of microscopic mechanisms behind the coherences’ thermal conductivity in Cu3BiS3, we calculate the spectral and cumulative mode-specific contributions to coherences’ conductivity at 300 and 400 K, respectively, as shown in Figure 3(d). In sharp contrast to the spectral populations’ conductivity of Cu3BiS3 in Figure 3(b), the majority of the coherences’ conductivity from the wavelike tunneling channel is carried by the phonons with almost full-frequency at both 300 and 400 K. This can be attributed to the small inter-band spacing within full-frequency region (due to dense phonon dispersions (see Figure S10 in the Supporting Information)) in conjunction with large linewidths (large scattering rates (see Figure 3(c)) or strong anharmonicity. Specifically, the spectral coherence conductivity exhibits a pronounced peak at 2.2 THz in the phase and at 1.8 THz in the Pnma phase. These peaks correspond to Einstein-like quasilocalized vibrational modes predominantly associated with Cu atoms (see Figure 3(b)). Within the middle and high-frequency domains, the contribution from spectral coherences’ conductivity is considerable, largely driven by strong anharmonic scattering alongside small interband spacing. Overall, for the Cu3BiS3,strong anharmonic effects combined with reduced phonon group velocities lead to quasilocalization in real space, thereby suppressing phonon-gas transport and fostering strong phonon coupling alongside significant diffuson contributions. Figures 4(a-f) show that the contributions to coherences’ thermal conductivity in Cu3BiS3 at 300 and 400 K, can be resolved in terms of the phonon energies, namely, and of two coupled phonons. At 300 K, similarly to the harmonic glasses 62, where quasi-degenerate phonon states dominate (see Figures 4(a-c)), the phonons contributing to along the three axis have relatively the same frequencies. For the pnma phase at 400 K, the two-dimensional density of states for the coherences’ thermal conductivity along the -aixs and -aixs (see Figures 4(d-e)) also shows couplings between quasi-degenerate states. In contrast, the the coherences’ thermal conductivity along the -aixs involves couplings between phonon modes with significantly different frequencies (see Figure 4(f)), driven by strong anharmonicity. As a result, the associated heat transport mechanism is fundamentally different from that in harmonic glasses.
3 Conclusion
In summary, we have comprehensively investigated the anharmonic lattice dynamics and thermal transport behavior in both ordered and partially disordered Cu3BiS3 using a combination of experimental measurements and advanced DFT-based theoretical simulations. To begin with, we experimentally synthesized and characterized both the ordered and liquid-like, partially Cu-occupied structures of Cu3BiS3, with the Pnma phase emerging upon temperature increase-an observation further supported by ab initio molecular dynamics simulations. Subsequently, we employ DFT-based self-consistent phonon theory with bubble-diagram corrections, combined with the Wigner transport equation, to capture the anharmonic lattice dynamics and uncover the microscopic mechanisms underlying the ultra-low thermal conductivity in Cu3BiS3. At low temperatures (2-100 K), the predicted thermal conductivity-accounting for 3ph, 4ph, isotope, 200 nm boundary, and point-defect scattering, accurately reproduce the experimental thermal conductivity. At 400 K, our results demonstrate that both the ordered phase and the partially Cu-occupied phase exhibit ultralow thermal conductivity, with values of 0.36 Wm-1K-1. Notably, our predicted thermal conductivity, incorporating both propagating and diffuson-like phonon contributions, well reproduces the experimental measurements of 0.34 Wm-1K-1, even for the partially Cu-occupied phase. In both the and phases, the ultralow thermal conductivity is mainly due to strong anharmonic scattering associated with Bi and Cu atoms. In particular, strong antibonding states between Cu and Se bonds give rise to the weak bonding of these Cu atoms, which leads to low-energy, Einstein-like vibrational modes that efficiently scatter heat-carrying acoustic phonons, thereby hindering their particle-like propagation. On the other hand, machine-learning-based Green–Kubo molecular dynamics simulations further confirm that ion hopping has a negligible impact on heat transport, even in the high-temperature, liquid-like phase. Therefore, despite the disordered nature of Cu3BiS3, our models—both machine learning and first-principles-based—provide reasonable estimates of thermal conductivity when using ordered structural representations. Our findings on Cu3BiS3 suggest that the lattice thermal conductivity () of amorphous or disordered compounds can, to a reasonable degree of accuracy, be approximated using calculations based on their ordered crystalline counterparts. This approach provides a practical pathway for evaluating thermal transport properties in complex systems where fully capturing structural disorder remains computationally challenging. Moreover, our study highlights the minimal impact of ionic mobility on thermal conductivity and offers valuable insights into the fundamental phonon transport mechanisms underlying the ultralow observed in both the ordered and liquid-like phases of Cu3BiS3.
J.Y. acknowledges the National Natural Science Foundation of China through grants No. 52072188 and No. 12204254, the Program for Science and Technology Innovation Team in Zhejiang through grant No. 2021R01004, as well as the Institute of High-pressure Physics of Ningbo University for its computational resources. X.S. acknowledges funding supported by National Youth Talent Project and the Fundamental Research Funds for the Central Universities (D5000250021), and the European Union’s Horizon 2020 research and innovation program under the Marie Sklodowska-Curie grant agreement No 101034329 and the WINNING Normandy Program supported by the Normandy Region. S.L. acknowledges the computation resources provided by the National Academic Infrastructure for Supercomputing in Sweden (NAISS) and the Swedish National Infrastructure for Computing (SNIC) at Sigma and Tetralith Clusters partially funded by the Swedish Research Council through grant agreements No. 2022-06725 and No. 2018-05973, as well as by the Vienna Scientific Cluster (VSC) in Austria.

### v2

@@
 This observation highlights the negligible impact of ionic mobility arising from partially occupied Cu sites on the thermal conductivity in diffuson-dominated thermal transport compounds.
-Our work not only sheds light on the minimal impact of ionic mobility on ultralow thermal conductivity in phase-change materials but also demonstrates that the Wigner transport equation accurately describes thermal transport behavior in partially occupied phases with diffuson-dominant thermal transport.
+Our work sheds light on the minimal impact of ionic mobility on ultralow thermal conductivity in phase-change materials.
+It demonstrates that the Wigner transport equation accurately describes thermal transport behavior in partially occupied phases with diffuson-dominant thermal transport.
 keywords:
-American Chemical Society, LaTeXThese authors contributed equally
-[…]
-Thayer School of Engineering, Dartmouth College, Hanover, New Hampshire, 03755, USA
-[…]
-[…]
-[…]
-Linköping University, Department of Physics, Chemistry, and Biology (IFM), Linköping, SE-58183, Sweden
-[…]
-State Key Laboratory of Superhard Materials, College of Physics, Jilin University, Changchun, 130012, China
-Keywords: thermoelectric material, Cu-based chalcogenide, lattice anharmonicity, wigner transport equation, self-consistent phonon theory
-[…]
-Some journals require a graphical entry for the Table of Contents. This should be laid out “print ready” so that the sizing of the text is correct.
-Inside the tocentry environment, the font used is Helvetica 8 pt, as required by Journal of the American Chemical Society.
-The surrounding frame is 9 cm by 3.5 cm, which is the maximum permitted for Journal of the American Chemical Society graphical table of content entries.
-The box will not resize if the content is too big: instead it will overflow the edge of the box.
-This box and the associated title will always be printed on a separate page at the end of the document.
-1 Introduction
-Ultra-low lattice thermal conductivity () in solid-state materials is crucial for advancing energy conversion technologies 1, 2, improving fuel cell performance 3, and optimizing various thermoelectric applications in the fields of thermal barrier coatings 4, waste heat recovery 5, 6.
-However, understanding and measuring lattice thermal conductivity is particularly challenging for materials undergoing phase transitions 7.
-These materials are widely available and have been extensively used in various fields, including thermoelectrics 8, solid-state memory 9, and switches 10.
+Ion mobility, ultra-low thermal conductivity, diffuson-dominated thermal transport, phase transition, machine-learning molecular dynamics simulation, Green-Kubo equation, Wigner transport equation1 Introduction
+Ultra-low lattice thermal conductivity () in solid-state materials is crucial for advancing energy conversion technologies [1, 2], improving fuel cell performance [3], and optimizing various thermoelectric applications in the fields of thermal barrier coatings [4], waste heat recovery [5, 6].
+However, understanding and measuring lattice thermal conductivity is particularly challenging for materials undergoing phase transitions [7].
+These materials are widely available and have been extensively used in various fields, including thermoelectrics [8], solid-state memory [9], and switches [10].
 Therefore, a fundamental understanding of the origins of ultra-low thermal conductivity through a comprehensive investigation of the relationships between transport phenomena, lattice dynamics, and structural configurations is of great significance.
-This understanding unveils critical insights and establishes foundational principles for the innovation and discovery of advanced material systems 11.
-Within the family of phase transition materials, Cu-based chalcogenides stand out due to their complex phase transition behaviors and outstanding thermoelectric properties 12, 13.
+This understanding unveils critical insights and establishes foundational principles for the innovation and discovery of advanced material systems [11].
+Within the family of phase transition materials, Cu-based chalcogenides stand out due to their complex phase transition behaviors and outstanding thermoelectric properties [12, 13].
 Cu-based chalcogenides often undergo complex lattice distortions and significant changes in thermal transport properties during phase transitions, making them a key platform for exploring new thermal conduction mechanisms.
 For instance, Liu et al.
-14 reported a significant reduction in the thermal conductivity of Cu2Se, attributed to its second-order phase transition, which led to a notable enhancement of the thermoelectric figure of merit () from 0.6 to 2.3.
-The recent synthesis of copper-rich wittichenite Cu3BiS3 15, 16, which exhibits intrinsically low lattice thermal conductivity, is particularly noteworthy.
+[14] reported a significant reduction in the thermal conductivity of Cu2Se, attributed to its second-order phase transition, which led to a notable enhancement of the thermoelectric figure of merit () from 0.6 to 2.3.
+The recent synthesis of copper-rich wittichenite Cu3BiS3 [15, 16, 17], which exhibits intrinsically low lattice thermal conductivity, is particularly noteworthy.
 This material provides a compelling platform to investigate structural features—such as atomic coordination and cationic networks—that influence thermal transport.
-Meanwhile, the present investigation on Cu3BiS3 clearly demonstrates that the 3-fold coordination of copper, combined with its strong anisotropic vibrations, serves as the driving force behind its ultra-low thermal conductivity 17.
-Despite extensive research on Cu3BiS3, its phase transition behavior remains unclear, and the mechanisms underlying its ultra-low thermal conductivity across different phases have yet to be fully elucidated.
-Historically, thermal transport in crystalline materials has been well described by the phonon-gas model, which treats phonons as quasiparticles propagating over long distances within the framework of the conventional phonon Boltzmann transport equation (BTE) 18, 19.
-Within the conventional BTE framework, point defects 20, anharmonicity 21, 22, and bond strength variations 23, 24 act as phonon scattering sources, reducing mean free paths and thereby lowering lattice thermal conductivity.
-However, certain complex crystalline materials exhibit ultra-low thermal conductivities that cannot be fully explained by the phonon-gas model 25.
-In such cases, phonon mean free paths are extremely short—comparable to or even smaller than interatomic distances—resulting in the loss of their quasi-particle-like nature 26.
+Meanwhile, Maji et al.
+demonstrate that the 3-fold coordination of copper, combined with its strong anisotropic vibrations, serves as the driving force behind the ultra-low thermal conductivity in Cu3BiS3 [17].
+Despite extensive researches on Cu3BiS3, its phase transition behavior remains unclear, and the mechanisms underlying its ultra-low thermal conductivity across different phases have yet to be fully elucidated.
+Historically, thermal transport in crystalline materials has been well described by the phonon-gas model, which treats phonons as quasiparticles propagating over long distances within the framework of the conventional Boltzmann transport equation (BTE) [18, 19].
+Within the conventional BTE framework, point defects [20], anharmonicity [21, 22], and bond strength variations [23, 24] act as phonon scattering sources, reducing mean free paths and thereby lowering lattice thermal conductivity.
+However, certain complex crystalline materials exhibit ultra-low thermal conductivities that cannot be fully explained by the phonon-gas model [25].
+The conventional BTE assumes well-separated phonon branches, and it therefore fails when this separation collapses, or when off-diagonal velocity terms become non-negligible—situations common in strongly anharmonic or disordered materials [26, 27].
 To address the limitations of the conventional BTE, Simoncelli et al.
-27 developed a unified theory of thermal transport by reformulating the Boltzmann Transport Equation within the Wigner formalism.
-Specifically, the Wigner transport equation 28 describes thermal transport via the phonon velocity operator, wherein the diagonal and off-diagonal elements capture both the particle-like propagation of phonons and the wave-like tunneling between eigenstates.
+[28] developed a unified theory of thermal transport by reformulating the Boltzmann Transport Equation within the Wigner formalism.
+Specifically, the Wigner transport equation [29] describes thermal transport via the phonon velocity operator, wherein the diagonal and off-diagonal elements (diffuson described by Allen and Feldman [26, 30]) capture both the particle-like propagation of phonons and the wave-like tunneling between eigenstates.
 The wave-like tunneling arises from the overlap of phonon modes with similar frequencies—substantially broadened by anharmonicity—thereby facilitating thermal energy transfer between modes.
-The unified theory of thermal transport has been successfully applied to address the underestimation of thermal conductivity within the BTE framework for strongly anharmonic systems, such as Cu12SbS29, La2Zr2O727, Cs2AgBiBr6 30.
-Another challenge in evaluating the thermal conductivity of phase-change materials is the potential occurrence of a partial-crystalline partial-liquid state (PCPLS), driven by the liquid-like migration of Cu ions, as observed in superionic systems such as Ag9GaSe631, AgCrSe232, and Li2S 33.
-It is also reported that the most representative -Cu2S, with space group and Pearson symbol cF208, exhibits a remarkably high degree of disorder, with 8 Cu ions distributed over 204 possible sites within the unit cell 34, 35.
+The Wigner transport equation has been successfully applied to address the underestimation of thermal conductivity within the BTE framework for strongly anharmonic systems, such as Cu12Sb4S13[31], La2Zr2O7[28], Cs2AgBiBr6 [32].
+Another challenge in evaluating the thermal conductivity of phase-change materials is the potential occurrence of a partial-crystalline partial-liquid state (PCPLS), driven by the liquid-like migration of Cu ions, as observed in superionic systems such as Ag9GaSe6[33], AgCrSe2[34], and Li2S [35].
+It is also reported that the most representative -Cu2S, with space group and Pearson symbol cF208, exhibits a remarkably high degree of disorder, with 8 Cu ions distributed over 204 possible sites within the unit cell [36, 37].
 The strikingly structural disorder suggests an exceptionally low activation barrier for Cu-ion migration, giving rise to liquid-like cationic dynamics.
-This remarkable ion mobility facilitates strong phonon scattering mechanisms, thereby reducing the specific heat via the suppression of transverse phonon modes 36.
+This remarkable ion mobility facilitates strong phonon scattering mechanisms, thereby reducing the specific heat via the suppression of transverse phonon modes [38].
 In contrast, Bernges et al.
-37 conducted the first direct comparison of thermal and ionic transport in nine Ag+ argyrodite compositions.
-Their results revealed that, despite the ionic conductivity varying by several orders of magnitude, the thermal conductivity remains largely constant, indicating that high ionic mobility is not essential for diffusion-mediated thermal transport.
+[39] conducted the first direct comparison of thermal and ionic transport in nine Ag+ argyrodite compositions.
+Their results revealed that, despite the ionic conductivity varying by several orders of magnitude, the thermal conductivity remains largely constant, indicating that high ionic mobility is not essential for diffuson-mediated thermal transport.
 Similarly, Ghata et al.
-13 observed that with increasing temperature, the ionic conductivity of Cu7PSe6 rises significantly, while the thermal conductivity remains nearly constant, suggesting a negligible influence of ionic transport on thermal conductivity.
+[13] observed that with increasing temperature, the ionic conductivity of Cu7PSe6 rises significantly, while the thermal conductivity remains nearly constant, suggesting a negligible influence of ionic transport on thermal conductivity.
 Therefore, a fundamental understanding of the thermal transport mechanism in phase-transition and partially occupied materials remains an open question and an urgent need.
@@
 We then combine first-principles calculations with machine-learning-based molecular dynamics simulations to accurately reproduce the partially Cu-atom occupied phase of Cu3BiS3 () at high temperatures.
-Subsequently, we applied the state-of-the-art Wigner transport equation, accounting for both propagative and diffuson contributions, to calculate the thermal conductivity of the ordered and phases of Cu3BiS3, achieving excellent agreement with experimental results.
-To investigate the impact of ion hopping on thermal transport at high temperatures in Cu3BiS3, we calculate the lattice thermal conductivity () using the machine-learning-based Green–Kubo equilibrium molecular dynamics (GK-EMD) method 38, which explicitly accounts for all contributions to the heat-flux operator 39, 40.
+Subsequently, we applied the state-of-the-art Wigner transport equation, accounting for both propagative and diffuson contributions, to calculate the thermal conductivity of the and phases of Cu3BiS3, achieving excellent agreement with experimental results.
+To investigate the impact of ion hopping on thermal transport at high temperatures in Cu3BiS3, we calculate the lattice thermal conductivity () using the machine-learning-based Green–Kubo equilibrium molecular dynamics (GK-EMD) method [40], which explicitly accounts for all contributions to the heat-flux operator [41, 42].
 We find that the predicted thermal conductivity accurately reproduces both the experimental data and the results obtained from the Wigner transport equation, indicating a negligible effect of ion mobility on thermal transport.
@@
 Our work not only highlights the negligible role of ionic mobility in thermal transport but also provides deeper insights into the microscopic mechanisms underlying the ultralow thermal conductivity observed in ordered and liquid-like Cu3BiS3 and related compounds.
-2 Results and discussion
-2.1 Experimental crystal structures
-Rietveld refinement of the powder X-ray diffraction (PXRD) data, carried out using the GSAS-II software package 41, 42, shows excellent agreement with the calculated patterns (R = 7.86%, R = 10.22%, and = 1.12), as illustrated in Figure 1(a).
-As evidenced by the well-defined PXRD patterns at 300 K, the compound is accurately indexed to the non-centrosymmetric space group , consistent with previously reported results 17, 43.
-Specifically, the structure comprises three Cu atoms, three S atoms, and one Bi atom, all positioned at Wyckoff 4a sites with a site multiplicity of 4.
-Each Bi atom is bonded to three S atoms, forming a [BiS3]3- trigonal pyramid that contains a lone pair of electrons.
-In contrast, each Cu atom is coordinated by three S atoms, forming a distorted trigonal planar arrangement.
-The unit cell parameters determined from Rietveld refinement of the phase-pure sample, namely, a = 7.6750(2) Å, b = 10.4083(3) Å, and c = 6.7060(2) Å, are in good agreement with density functional theory (DFT) calculations performed using the PBEsol functional 44 (a = 7.4618(8) Å, b = 10.4705(9) Å, c = 6.3209(5) Å) as well as with previously reported results 16, 17.
-Upon further increasing the temperature, we observe a notable change in the XRD pattern (see Figure S1 in the Supporting Information).
-Based on the structural refinments of the high-T PXRD pattern, we identify an experimental phase transition from the orthorhombic space group to the orthorhombic space group is observed, as illustrated in Figure 1(b).
-This finding diverges from the previously proposed space group for CuBiS2 reported by Wu et al.in Ref.16.
-Although our results differ from those reported by Wu et al., it is noteworthy that the isostructural space Cu3SbS3 has been experimentally observed to undergo a phase transition from the (-) phase to the (-) phase 45, 46.
-This observation aligns with our findings, supporting the validity of the observed phase transition in Cu3BiS3.
-Interestingly, the distribution of disordered Cu+ ions within the three-dimensional framework of [BiS3]3- units has been elucidated in the phase, matching that observed in -Cu3SbS3 47.
-Specifically, single-crystal data reveal that Cu+ ions exhibit disorder, being partially distributed over five distinct sites, resulting in either trigonal-planar or tetrahedral coordination.
-To investigate this phenomenon, we conducted 20 ps ab initio molecular dynamics (AIMD) simulations at 500 K (see Figure 1(c) and Figure S2 in the Supporting Information).
+2 Experimental crystal structures
+Rietveld refinement carried out using the Fullprof software package [43, 44], leads to calculated patterns in excellent agreement with the powder X-ray diffraction (PXRD) data, as illustrated in Figure 1.
+As evidenced by the well-defined refinement at 300 K (Rp = 7.96, Rwp = 10.3, Rexp = 9.13 and = 1.28), the PXRD pattern is accurately indexed to the non-centrosymmetric space group , consistent with previously reported results [17, 45, 46].
+Specifically, the crystal structure is characterized by Cu, Bi and S atoms located on three, one and three differents crystallographic sites, respectively, of general position, i.e.
+Wyckoff 4a.
+Each Bi atom is bonded to three S atoms at distances ranging between 2.565 Å and 2.614 Å (average Bi-S distance of 2.594 Å), forming a [BiS3]3- trigonal pyramid.
+Note that the lone pair (LP) of electrons of Bi3+ cation very probably constitutes the missing apex of a tetrahedron [BiS3LP]3-.
+In contrast, each Cu atom is coordinated by three S atoms at distances ranging between 2.178 Å and 2.381 Å (average Cu-S distance of 2.285 Å), forming a distorted trigonal planar arrangement.
+The unit cell parameters determined from Rietveld refinement of the phase-pure sample, namely, a = 7.6769(2) Å, b = 10.4107(3) Å, and c = 6.7076(2) Å, are in agreement with density functional theory (DFT) calculations performed using the PBEsol functional [47] (a = 7.4618(8) Å, b = 10.4705(9) Å, c = 6.3209(5) Å) as well as with previously reported results [16, 17].
+Upon further increasing the temperature, we observe a notable change in the XRD pattern (see Figure S1 in the Supporting Information) confirming previous results reported by Makovicky highlighting that Cu3BiS3 exhibits a structural transition from a RT orthorhombic () structure to an intermediate modulated structure at 391.5 K and, above 463 K, a transition to a high temperature (HT) polymorph () [46].
+Based on the structural refinments of the high-T PXRD patterns, we confirm an experimental phase transition from the orthorhombic space group to the orthorhombic space group , as illustrated in Figure 1, even if the space group is not excluded [46].
+This space group change from at 300 K to at 500 K is related to the extinction of the (hk0) reflections for which h = 2n+1, as exemplified with the (120) and (310) reflections observed at 20.6° and 36.1° in 2 on the PXRD pattern recorded at 300 K and absent at 500 K.
+In addition, a broad peak observed near 2 13.2° does not match the known reflections of the Cu3BiS3 phase and is significantly wider than the main phase peaks.
+This suggests that it may arise from the sample environment or from a poorly crystallized impurity phase that could not be reliably identified.
+On the other hand, we have observed that CuBiS2 exists in the crystal system based on the structural data reported by Ref.
+[48].
+To rule out the possibility of CuBiS2 as a minority or mis-indexed phase, we simulated powder-diffraction patterns for the three most plausible structures at the Cu, K wavelength: (i) Cu3BiS3 in and , the two space groups obtained from our Rietveld refinements, and (ii) CuBiS2 in , using the atomic coordinates reported by Portheine and Nowacki [48] (see Figure S2 in the Supporting Information).
+The calculated pattern for CuBiS2 displays an intense reflection at 2 12°, a feature that is completely absent in our experimental data, where the first detectable peak appears above 15°.
+Additional reflections predicted for CuBiS2 (e.g., 2 18.3°, 26.9°, 31.7°) are also missing or mismatched in both position and intensity.
+These systematic discrepancies unambiguously exclude CuBiS2 as a constituent phase in the investigated sample.
+To quantify the competition between ordered and liquid-like states from first principles, we have constructed a Cu–Bi–S phase diagram based on the convex-hull of DFT formation enthalpies (see Figure S3(a) in the Supporting Information).
+The diagram encompasses all computed compositions and explicitly resolves the two Cu3BiS3 polymorphs: the orthorhombic phase and the ordered variant.
+Notably, the structure lies 3.16 meV/atom below , i.e.
+= –3.16 meV/atom, confirming its thermodynamic preference.
+Interestingly, the distribution of Cu+ ions within the framework of [BiS3]3- units has been elucidated in the phase (see Figure S3(b) in the Supporting Information), matching that observed in -Cu3SbS3 [49].
+Specifically, single-crystal data recorded at 493 K reveal that Cu+ ions exhibit disorder (see Figure S3(c) in the Supporting Information), being distributed over five distinct sites each being partially occupied, resulting in either trigonal-planar or tetrahedral coordination.
+To investigate this phenomenon, we conducted 20 ps ab initio molecular dynamics (AIMD) simulations at 400 K and 500 K (see Figure 2(a) and Figure S3(d-e) in the Supporting Information).
 Notably, the simulations successfully captured the diverse sites occupied by Cu+ ions, confirming their mobility and aligning closely with our experimental observations.
 To elucidate the chemical environment of Cu+ ions, we performed electronic structure calculations on the phase, examining both bonding and antibonding states to gain deeper insights into its bonding characteristics.
-In the phase, various Cu–S bonds exhibit distinct bonding characteristics (see Figure 1(d)); nevertheless, these bonds generally share common features.
-All Cu–Se bonds exhibit occupied antibonding states below the Fermi level, contributing approximately 45% to the overall antibonding character.
-Previous studies have attributed this behavior to the proximity of the Cu d-orbitals to the energy levels of the S p-orbitals, leading to strong p–d hybridization 48.
-Such robust hybridization pushes the occupied antibonding states closer to the Fermi level, weakening the chemical bonds and promoting metavalent-like behavior 24, 49.
-The combination of low integrated electron population and substantial antibonding contributions highlights the intrinsically weak bonding environment of Cu+ within the structure, thus facilitating rapid Cu+ ion transport 50.
+In the phase, various Cu–S bonds exhibit distinct bonding characteristics (see Figure 2(b)); nevertheless, these bonds generally share common features.
+All Cu–S bonds exhibit occupied antibonding states below the Fermi level, contributing approximately 45% to the overall antibonding character.
+Previous studies have attributed this behavior to the proximity of the Cu d-orbitals to the energy levels of the S p-orbitals, leading to strong p–d hybridization [50].
+Such robust hybridization pushes the occupied antibonding states closer to the Fermi level, weakening the chemical bonds and promoting metavalent-like behavior [24, 51].
+The combination of low integrated electron population and substantial antibonding contributions highlights the intrinsically weak bonding environment of Cu+ within the structure, thus facilitating rapid Cu+ ion transport [52].
 Furthermore, the significant presence of occupied antibonding states below the Fermi level notably enhances lattice anharmonicity, substantially influencing thermal transport, as discussed in subsequent sections.
 Also, such a weak bonding leads to larger mean square displacement (MSD) of Cu atoms compared to the other atoms (see Figure S4 in the Supporting Information).
-This is consistent with the experimental observation of the atomic displacement parameters of Cu3BiS3 atoms 17.
-2.2 Diffuson-dominanted thermal transport
-To accurately capture the finite-temperature lattice dynamics, the anharmonic phonon frequencies of crystalline Cu3BiS3 are calculated using self-consistent phonon theory with bubble correction (SCPB) 30, 51 (see Figures 2(a) and 2(b)).
-Interestingly, from Figures 2(a) and 2(b), we find that in both the and Pnma phases, low-frequency phonons exhibit softening with increasing temperature, in significant contrast to previously reported results for other crystalline compounds, such as BaZrO3 26, TlInTe2 52, and YbFe4Sb12 53.
-Note that in strongly anharmonic materials, such as CsPbBr351, KNbO3and NaNbO354, the negative energy shift from the bubble diagram 51 is significant and imposes a non-nigligible impact on both the theoretical phase transition temperature and the phonon linewidth.
-Specifically, considering only the energy shift from quartic anharmonicity leads to low-frequency phonon hardening in both the and Pnma phases of Cu3BiS3 (see Figure S4 in the Supporting Information).
+This is consistent with the experimental observation of the atomic displacement parameters of Cu atoms reported in Cu3BiS3 [17].
+3 Diffuson-dominated thermal transport
+To accurately capture the finite-temperature lattice dynamics, the anharmonic phonon frequencies of crystalline Cu3BiS3 are calculated using self-consistent phonon theory with bubble correction (SCPB) [32, 53] (see Figures 3(a) and 3(b)).
+Interestingly, from Figures 3(a) and 3(b), we find that in both the and phases, low-frequency phonons exhibit softening with increasing temperature, in significant contrast to previously reported results for other crystalline compounds, such as BaZrO3 [27], TlInTe2 [54], and YbFe4Sb12 [55].
+Note that in strongly anharmonic materials, such as CsPbBr3 [53], KNbO3 and NaNbO3 [56], the negative energy shift from the bubble diagram [53] is significant and imposes a non-negligible impact on both the theoretical phase transition temperature and the phonon linewidth.
+Specifically, considering only the energy shift from quartic anharmonicity leads to low-frequency phonon hardening in both the and phases of Cu3BiS3 (see Figure S5 in the Supporting Information).
 Therefore, incorporating the full explicit anharmonic vibrational effects, including both cubic and quartic anharmonicity, enables a more accurate description of phonon softening, which significantly influences the lattice thermal conductivity.
-With the finite-temperature phonons obtained, we proceed to calculate the lattice thermal conductivity of Cu3BiS3 using the Wigner transport equation 27, 28, accounting for both phonon population and coherence contributions, as shown in Figure 2(c).
-Our results show that the thermal conductivity of the phase at 400 K reaches 0.36 Wm-1K-1, in good agreement with the experimental value of 0.34 Wm-1K-1 17, thereby validating our calculations.
-More specifically, at 400 K, the coherence contribution to the total thermal conductivity is 0.22 Wm-1K-1, accounting for 6̃1% of the total thermal conductivity, highlighting the dominant role of phonon wave-like tunneling and strong lattice anharmonicity 55.
-We next continue to calculate the thermal conductivity of the ordered Pnma phase of Cu3BiS3 above 400 K using the Wigner transport equation 27, 28(see in Figure 2(c)).
+With the finite-temperature phonons obtained, we proceed to calculate the lattice thermal conductivity of Cu3BiS3 using the Wigner transport equation [28, 29], accounting for both phonon population and coherence contributions, as shown in Figure 3(c).
+Our results show that the thermal conductivity of the phase at 400 K reaches 0.36 Wm-1K-1, in good agreement with the experimental value of 0.34 Wm-1K-1 [17], thereby validating our calculations.
+More specifically, at 400 K, the coherence contribution to the total thermal conductivity is 0.22 Wm-1K-1, accounting for 6̃1% of the total thermal conductivity, highlighting the dominant role of phonon wave-like tunneling and strong lattice anharmonicity [57].
+We next continue to calculate the thermal conductivity of the phase of Cu3BiS3 above 400 K using the Wigner transport equation [28, 29](see in Figure 3(c)).
 Our theoretical model predicts an ultra-low total thermal conductivity of 0.36 Wm-1K-1, comprising 0.16 Wm-1K-1 from the population contribution and 0.20 Wm-1K-1 from the coherence contribution.
-Surprisingly, the predicted thermal conductivity of the ordered Pnma phase of Cu3BiS3 agrees well with the experimentally observed value of 0.34 Wm-1K-1 at 400 K for the partially occupied Pnma phase.
-Furthermore, as temperature increases, the theoretical predictions remain in good agreement with experimental measurements for partial-occupied Cu3BiS3 (Pnma space group), as shown in Figure 2(c).
-The cross-phase thermal transport in Cu3BiS3 remains unexplored 16, motivating us to compare the predicted thermal conductivities of both the and Pnma phases with experimental values, as shown in the inset of Figure 2(c).
+Surprisingly, the predicted thermal conductivity of the phase of Cu3BiS3 agrees well with the experimentally observed value of 0.34 Wm-1K-1 at 400 K for the partially occupied phase.
+Furthermore, as temperature increases, the theoretical predictions remain in good agreement with experimental measurements for partial-occupied Cu3BiS3 ( space group), as shown in Figure 3(c).
+The cross-phase thermal transport in Cu3BiS3 remains unexplored [16], motivating us to compare the predicted thermal conductivities of both the and phases with experimental values, as shown in the inset of Figure 3(c).
 Our results show that the thermal conductivity of the phase increases sharply above 400 K, deviating from the experimental values.
-In contrast, the Pnma phase shows a continuous decrease in thermal conductivity above 400 K, showing a good agreement with experimental results at high temperatures.
-These results further confirm the temperature-induced phase transition from to Pnma and provide a clear explanation for the smooth change in thermal conductivity at the transition point.
-Considering the liquid-like behavior of the Pnma phase at high temperatures, we calculate using the Green-Kubo equilibrium molecular dynamics (GK-EMD) method 38 (see Figure S5 in the Supporting Information), which accounts for all terms in the heat flux operator 39, 40.
-As shown in Figure 2(c), the total thermal conductivity predicted by the GK-EMD method at 400 K is in excellent agreement with that obtained from the Wigner transport equation 27, 28.
+In contrast, the phase shows a continuous decrease in thermal conductivity above 400 K, showing a good agreement with experimental results at high temperatures.
+These results further confirm the temperature-induced phase transition from to and provide a clear explanation for the smooth change in thermal conductivity at the transition point.
+It is worth noting that the Wigner transport equation is limited by its reliance on the anharmonic heat flux contributions [29], and higher-order phonon scattering.
+Moreover, it cannot capture liquid-like ion diffuson [58], such as the motion of ions in disordered or partially liquid environments.
+In contrast, molecular dynamics methods naturally include these effects, making them more suitable for such scenarios.
+Considering the liquid-like behavior of the phase at high temperatures, we calculate using the Green-Kubo equilibrium molecular dynamics (GK-EMD) method [40] (see Figure S6 in the Supporting Information), which accounts for all terms in the heat flux operator [41, 42].
+As shown in Figure 3(c), the total thermal conductivity predicted by the GK-EMD method at 400 K is in excellent agreement with that obtained from the Wigner transport equation [28, 29].
 Even at 500 K, the discrepancy between the two methods remains negligible, with a difference of only 0.01 Wm-1K-1.
-Based on the above observation, and considering that the thermal conductivity from the Wigner transport equation 27, 28 is calculated for the ordered Pnma phase—i.e., in the absence of ion hopping—we conclude that ion hopping has minimal impact on thermal transport in liquid-like Cu3BiS3.
-The negligible effect of ion hopping on thermal transport can be attributed to the fact that ion-hopping behavior resembles rattling-like motion observed in certain crystalline materials 13, 56, which induces strong anharmonicity.
+Based on the above observation, and considering that the thermal conductivity from the Wigner transport equation [28, 29] is calculated for the phase—i.e., in the absence of ion hopping—we conclude that ion hopping has minimal impact on thermal transport in liquid-like Cu3BiS3.
+The negligible effect of ion hopping on thermal transport can be attributed to the fact that ion-hopping behavior resembles rattling-like motion observed in certain crystalline materials [13, 59], which induces strong anharmonicity.
 This, in turn, suppresses phonon transport while enhancing the coherence contribution.
@@
 To further validate our theoretically predicted lattice thermal conductivity of crystalline Cu3BiS3, we compared the calculated values with experimental measurements in the temperature range of 2 to 100 K (see Figure 3(d)).
-Clearly, when boundary scattering at low temperatures is not considered 18, the predicted thermal conductivity—accounting only for three-phonon, four-phonon, and phonon-isotope scattering—deviates significantly from the experimental values.
-To accurately reproduce the experimental thermal conductivity 57, 58, effective grain sizes of 50, 100, and 200 nm are introduced, which significantly suppress the thermal conductivity and bring the predictions closer to the experimental results.
+Clearly, when boundary scattering at low temperatures is not considered [18], the predicted thermal conductivity—accounting only for three-phonon, four-phonon, and phonon-isotope scattering—deviates significantly from the experimental values.
+To accurately reproduce the experimental thermal conductivity [60, 61], effective grain sizes of 50, 100, and 200 nm are introduced, which significantly suppress the thermal conductivity and bring the predictions closer to the experimental results.
 On top of the 200 nm boundary scattering model, the inclusion of additional point-defect scattering leads to improved agreement between theory and experiment, as depicted in Figure 3(d).
@@
 Meanwhile, our computational results closely reproduce the isobaric heat capacity data reported by Wu et al.
-prior to the observed phase transition 16.
-2.3 Microscopic mechanisms of thermal transport
+prior to the observed phase transition [16].
+4 Microscopic mechanisms of thermal transport
 Next, to reveal the microscopic mechanisms of thermal transport in crystalline Cu3BiS3 with the and phases, we calculate the spectral and cumulative propagating thermal conductivity at 300 K and 400 K, respectively.
@@
 In contrast, S atoms primarily participate in the high-frequency optical modes above 5.0 THz, while Bi atoms exhibit relatively low participation and are mainly involved in a limited number of low-frequency optical modes between 1.5 and 2.0 THz (see Figure S9 in the Supporting Information).
-The fully Cu- and Bi-occupied low-frequency optical modes lead to strong phonon scattering rates, thereby suppressing thermal transport in crystalline Cu3BiS3, a phenomenon similarly observed in CsCu4Se3 55.
-Importantly, Cu atoms give rise to low-energy, Einstein-like quasilocalized vibrational modes 56, particularly within the 2–4 THz frequency range.
+The fully Cu- and Bi-occupied low-frequency optical modes lead to strong phonon scattering rates, thereby suppressing thermal transport in crystalline Cu3BiS3, a phenomenon similarly observed in CsCu4Se3 [57].
+Importantly, Cu atoms give rise to low-energy, Einstein-like quasilocalized vibrational modes [59], particularly within the 2–4 THz frequency range.
 These Cu-dominated quasilocalized vibrational phonon modes act as strong phonon scatterers, significantly interacting with and disrupting the propagation of heat-carrying acoustic phonons.
-This phenomenon has also been reported in cubic SrTiO359, the double perovskite Cs2AgBiBr630, and copper-based chalcogenides such as o-CsCu5S324.
+This phenomenon has also been reported in cubic SrTiO3[62], the double perovskite Cs2AgBiBr6[32], and copper-based chalcogenides such as o-CsCu5S3[24].
 As illustrated in Figure 4(c), we further calculate the phonon lifetimes for the phase at 300 K and the phase at 400 K, respectively.
-We find that the majority of phonon modes exhibit lifetimes shorter than the Wigner limit 27, 28, considering both three-phonon (3ph) and four-phonon (4ph) scattering processes, indicating strong anharmonicity in the system.
-For the phase, we observe a dip in phonon lifetimes for phonons around 0.75 THz, which are dominated by Bi atoms, highlighting a strong scattering source that contributes to the suppression of thermal conductivity 26, 30.
+We find that the majority of phonon modes exhibit lifetimes shorter than the Wigner limit [28, 29], considering both three-phonon (3ph) and four-phonon (4ph) scattering processes, indicating strong anharmonicity in the system.
+For the phase, we observe a dip in phonon lifetimes for phonons around 0.75 THz, which are dominated by Bi atoms, highlighting a strong scattering source that contributes to the suppression of thermal conductivity [27, 32].
 Similarly, the flattened, localized vibrational mode, primarily associated with Cu atoms near 2.2 THz, also exhibits an extremely short lifetime and pronounced anharmonic scattering (see Figure 4(c) and Figure S10 in the Supporting Information).
@@
 In the phase, Cu atoms play a similar role, as the Cu-dominated mode near 1.8 THz exhibits a significant drop in phonon lifetime (see Figure 4(c) and Figure S10 in the Supporting Information), corresponding to the dip observed in the spectral thermal conductivity (see Figure 4(a)).
-Although strong anharmonicity and the low phonon lifetimes of quasilocalized modes suppress particle-like phonon propagation 25, they enhance the wave-like phonon diffuson contribution (see Figure 3(d)).
-Theoretically, the presence of dense, quasilocalized phonon modes with similar energies and strong anharmonicity can facilitate significant heat exchange between phonon modes in energy space 60, 55, thereby promoting diffusion-like transport 61.
+Although strong anharmonicity and the low phonon lifetimes of quasilocalized modes suppress particle-like phonon propagation [25], they enhance the wave-like phonon diffuson contribution (see Figure 4(d)).
+Theoretically, the presence of dense, quasilocalized phonon modes with similar energies and strong anharmonicity can facilitate significant heat exchange between phonon modes in energy space [63, 57], thereby promoting diffuson-like transport [64].
 To gain a better understanding of microscopic mechanisms behind the coherences’ thermal conductivity in Cu3BiS3, we calculate the spectral and cumulative mode-specific contributions to coherences’ conductivity at 300 and 400 K, respectively, as shown in Figure 4(d).
@@
 This can be attributed to the small inter-band spacing within full-frequency region (due to dense phonon dispersions (see Figure S11 in the Supporting Information)) in conjunction with large linewidths (large scattering rates (see Figure 4(c)) or strong anharmonicity.
-Specifically, the spectral coherence conductivity exhibits a pronounced peak at 2.2 THz in the phase and at 1.8 THz in the Pnma phase.
+Specifically, the spectral coherence conductivity exhibits a pronounced peak at 2.2 THz in the phase and at 1.8 THz in the phase.
 These peaks correspond to Einstein-like quasilocalized vibrational modes predominantly associated with Cu atoms (see Figure 4(b)).
@@
 Figures 5(a-f) show that the contributions to coherences’ thermal conductivity in Cu3BiS3 at 300 and 400 K, can be resolved in terms of the phonon energies, namely, and of two coupled phonons.
-At 300 K, similarly to the harmonic glasses 62, where quasi-degenerate phonon states dominate (see Figures 4(a-c)), the phonons contributing to along the three axis have relatively the same frequencies.
-For the pnma phase at 400 K, the two-dimensional density of states for the coherences’ thermal conductivity along the -aixs and -aixs (see Figures 4(d-e)) also shows couplings between quasi-degenerate states.
+At 300 K, similarly to the harmonic glasses [65], where quasi-degenerate phonon states dominate (see Figures 5(a-c)), the phonons contributing to along the three axe have relatively the same frequencies.
+For the phase at 400 K, the two-dimensional density of states for the coherences’ thermal conductivity along the -aixs and -aixs (see Figures 5(d-e)) also shows couplings between quasi-degenerate states.
 In contrast, the the coherences’ thermal conductivity along the -aixs involves couplings between phonon modes with significantly different frequencies (see Figure 5(f)), driven by strong anharmonicity.
@@
 In summary, we have comprehensively investigated the anharmonic lattice dynamics and thermal transport behavior in both ordered and partially disordered Cu3BiS3 using a combination of experimental measurements and advanced DFT-based theoretical simulations.
-To begin with, we experimentally synthesized and characterized both the ordered and liquid-like, partially Cu-occupied structures of Cu3BiS3, with the Pnma phase emerging upon temperature increase-an observation further supported by ab initio molecular dynamics simulations.
+To begin with, we experimentally synthesized and characterized both the ordered structure stable at room temperature and liquid-like structure occurring above 400 K of Cu3BiS3.
+The latter disordered structure crystallizing in the space group, is characterized by statistical distribution of Cu+ ions on several crystallographic sites and emerges upon temperature increase - an observation further supported by ab initio molecular dynamics simulations.
 Subsequently, we employ DFT-based self-consistent phonon theory with bubble-diagram corrections, combined with the Wigner transport equation, to capture the anharmonic lattice dynamics and uncover the microscopic mechanisms underlying the ultra-low thermal conductivity in Cu3BiS3.
 At low temperatures (2-100 K), the predicted thermal conductivity-accounting for 3ph, 4ph, isotope, 200 nm boundary, and point-defect scattering, accurately reproduce the experimental thermal conductivity.
-At 400 K, our results demonstrate that both the ordered phase and the partially Cu-occupied phase exhibit ultralow thermal conductivity, with values of 0.36 Wm-1K-1.
+At 400 K, our results demonstrate that both the ordered phase and disordered phase exhibit ultralow thermal conductivity, with values of 0.36 Wm-1K-1.
 Notably, our predicted thermal conductivity, incorporating both propagating and diffuson-like phonon contributions, well reproduces the experimental measurements of 0.34 Wm-1K-1, even for the partially Cu-occupied phase.
 In both the and phases, the ultralow thermal conductivity is mainly due to strong anharmonic scattering associated with Bi and Cu atoms.
-In particular, strong antibonding states between Cu and Se bonds give rise to the weak bonding of these Cu atoms, which leads to low-energy, Einstein-like vibrational modes that efficiently scatter heat-carrying acoustic phonons, thereby hindering their particle-like propagation.
+In particular, large anisotropic vibrations of Cu in triangular coordination leads to low-energy, Einstein-like vibrational modes that efficiently scatter heat-carrying acoustic phonons, thereby hindering their particle-like propagation.
 On the other hand, machine-learning-based Green–Kubo molecular dynamics simulations further confirm that ion hopping has a negligible impact on heat transport, even in the high-temperature, liquid-like phase.
@@
 Moreover, our study highlights the minimal impact of ionic mobility on thermal conductivity and offers valuable insights into the fundamental phonon transport mechanisms underlying the ultralow observed in both the ordered and liquid-like phases of Cu3BiS3.
+Acknowledgements
 J.Y.
@@
 acknowledges funding supported by National Youth Talent Project and the Fundamental Research Funds for the Central Universities (D5000250021), and the European Union’s Horizon 2020 research and innovation program under the Marie Sklodowska-Curie grant agreement No 101034329 and the WINNING Normandy Program supported by the Normandy Region.
+X.S, K.M, and E.G gratefully thank Jean-François Lohier, Fabien Veillon, Christelle Bilot, and Jerôme Lecourt for technical support.
 S.L.
-acknowledges the computation resources provided by the National Academic Infrastructure for Supercomputing in Sweden (NAISS) and the Swedish National Infrastructure for Computing (SNIC) at Sigma and Tetralith Clusters partially funded by the Swedish Research Council through grant agreements No.
-2022-06725 and No.
-2018-05973, as well as by the Vienna Scientific Cluster (VSC) in Austria.
+acknowledges the computation resources provided by the National Academic Infrastructure for Supercomputing in Sweden (NAISS) – at supercomputer centers NSC and PDC, partially funded by the Swedish Research Council through grant agreement no.
+2022-06725 – and by the Vienna Scientific Cluster (VSC) in Austria.
+Conflict of interest
+The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.